I have a word problem that I'm trying to help my child with. I have the answer, but can't visualize how to get to it. I need help please. Here goes:
Stu had quite a collection of baseball cards. One day he decided to give half of them to his brother Seth. Then he gave one-third of what he had left to a friend. Then he gave half of what was left to another friend. Stu now has 50 baseball cards. How many did he start with? It seems that I should work backwards, but I'll be darned if I can figure out how. The answer is 300. This is a math enrichment problem in his homeschool co-op binder from Evan-Moor 6th grade. Any help would be appreciated as to how to do and explain why we are doing what we are doing. Lisa T.

Answer by mananth(16946)

Stu had quite a collection of baseball cards. One day he decided to give half of them to his brother Seth. Then he gave one-third of what he had left to a friend. Then he gave half of what was left to another friend. Stu now has 50 baseball cards. How many did he start with? It seems that I should work
let Stu have x cards
he gave half to seth
so he had x/2 cards
He had x/2 he gave 1/3 of x/2 to a friend
=x/6.
Now he had x/2 - x/6 = x/3 with him balance
1/2 of x/3 he gave to another friend =x/6
Balance he had was x/3 - x/6 = x/6
x/6 = 50
multiply by 6
x= 300